OpenStudy (anonymous):

Please help! A developer borrows $150,000 at 6.5% interest, compounded monthly and agrees to pay off the loan in 5 years. How much interest will she have paid?

OpenStudy (anonymous):

help ppl ? not sure how to do it

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) A = 150000(1+0.065/12)^(12*5) A = 207,422.5986 A = 207,422.60 So after 5 years, $207,422.60 is paid back in total. The amount in interest paid back is then 207,422.60 - 150,000 = 57,422.6 dollars

OpenStudy (anonymous):

thank you so muuchhh! would you like to be my friend:)

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (anonymous):

:)